# Learning to Generate Parameters of ConvNets for Unseen Image Data

### v1

@@ III-D Context-Aware Parameter Prediction
After obtaining the sketches for all training datasets, we will feed them into the context-aware parameter prediction module, as shown in Fig. 2. In the followings, we will introduce this module in detail.
III-D1 Capturing Contextual Parameter Relations via Adaptive Hyper Recurrent Units
First, since we aim to establish a hypernetwork that can capture the hyper-mapping between datasets and their corresponding network parameters, the hypernetwork should be dataset-agnostic. Second, as the input of a neural network would sequentially pass forward its layers, the parameters of different layers should be not independent. Neglecting the contextual relations among parameters of different layers may result in sub-optimal solutions.
In light of this, we construct Adaptive Hyper Recurrent Units (AHRU), specifically designed for parameter prediction of ConvNets. AHRU enhances the capabilities of the gated recurrent unit by adapting to different datasets, while also acting as the hyper recurrent unit to capture context-aware parameter relations during the parameter generation process. As shown in Fig. 3, we set the dataset sketch embeddings as the input of the hyper recurrent unit:
$\mathbf{d}_{0}^{(i)}=\mathbf{s}_{i},\quad\mathbf{s}_{i}\in\{\mathbf{s}_{i}\}_{i=1}^{\mathcal{N}}.$ (4)
By this means, we can provide dataset information for predicting the corresponding parameter representations. Then, to predict parameter representation $\mathbf{a}_{t}^{(i)}$ of $t$-th layer, we exploit the context data information $\mathbf{d}_{t-1}^{(i)}$ as the hidden state and the predicted parameter representation $\mathbf{a}_{t-1}^{(i)}$ as the input, enabling the information of shallower layer parameters memorized in AHRU to help the prediction of parameters in the deeper layer. By this way, the dependency relations among parameters of different layers can be well captured.
@@ III-D1 Capturing Contextual Parameter Relations via Adaptive Hyper Recurrent Units
The following comes a formal description:
@@
where $\mathbf{d}_{t}^{(i)}$ could transfer dataset-related context information for $i$-th dataset. $\mathbf{a}_{t}^{(i)}$ is the predicted parameter representation. $\mathbf{W}_{r},\mathbf{W}_{z},\mathbf{W}_{h},\mathbf{W}_{o}$ are learnable parameters shared by different datasets. In AHRU, the reset gate $\mathbf{r}_{t}^{(i)}$ decides how much data information in hidden state $\mathbf{d}_{t-1}^{(i)}$ needs to be reset. $\tilde{\mathbf{d}}_{t}^{(i)}$ is a new memory, which absorbs the information of $\mathbf{d}_{t-1}^{(i)}$ and $\mathbf{a}_{t-1}^{(i)}$. $\mathbf{z}_{t}$ is an update gate, which regulates how much information in $\tilde{\mathbf{d}}_{t}^{(i)}$ to update and how much information in $\mathbf{d}_{t-1}^{(i)}$ to forget.
Note that, AHRU is designed based on the GRU-based architecture [52]. Differently, our AHRU is applied to capture the dependencies among parameters of different layers, which can adapt to various datasets. Moreover, conventional GRU utilizes the randomly initialized hidden state to carry the long-term memory. In each recurrent step, it needs to receive external input (e.g., word) to update the hidden state and capture the contextual information. Instead of that, we set each dataset sketch embedding as input and follow a self-loop structure to incorporate the predicted structure parameter representation from the previous layer. This approach allows us to leverage the dataset-specific information and meanwhile enable the information of shallower layer parameters to help the prediction of parameters in the deeper layer. By this way, the dependencies of parameters of different layers can be well captured.
@@ III-D2 Initial Residual Connection
To ensure that the final context-aware output contains at least a fraction of the initial dataset information, we additionally implement an initial residual connection between dataset sketch embedding $\mathbf{s}_{i}$ and $\mathbf{a}_{t}$ as:
$\hat{\mathbf{a}}_{t}^{(i)}=\mathbf{a}_{t}^{(i)}\times(1-\eta)+\mathbf{s}_{i}^{(i)}\times\eta,\quad\hat{\mathbf{a}}_{t}^{(i)}\in\mathbb{R}^{m},$ (5)
where $\eta$ is the hyperparameter. After obtaining $\hat{\mathbf{a}}_{t}$, we put $\hat{\mathbf{a}}_{t}$ into the weight generator to generate parameters of the $t$-th layer of the target network $\Omega$.
@@ III-D3 Weight Generator
Since the target network $\Omega$ usually has different sizes and dimensions in different layers, we construct the weight generator $g_{\psi_{t}}$ for each layer $t$ to transform $\hat{\mathbf{a}}_{t}^{(i)}$ of a fixed dimension to network parameter tensor $\mathbf{w}_{t}^{(i)}$ with variable dimensions. Here $g_{\psi_{t}}$ denotes the weight generator of the $t$-th layer, $\psi_{t}$ is the learnable parameters of $g_{\psi_{t}}$, $\mathbf{w}_{t}^{(i)}$ is the predicted parameter of the $t$-th layer in $\Omega$ for the $i$-th dataset. We can derive the predicted parameter of the $t$-th layer as:
@@
where $g_{\psi_{t}}$ consists of one linear layer and two $1\times 1$ convolutional layers. Fig. 4 shows the architecture of the weight generator. The weight generator takes the vector $\hat{\mathbf{a}}_{t}^{(i)}$ with the dimension of $m$ as input, and outputs a tensor with the size of $C_{out}\times C_{in}\times k^{2}$ as the parameters of the $t$-th convolutional layer.
@@ III-E Optimization of Our Framework
In this section, we introduce how to optimize our PudNet. In contrast to traditional classification tasks where training data and testing data have the identical label space. In our task, the label spaces between training and testing can be different, even not overlapped. Thus, training a classification head on training data can not be used to predict labels of testing data. Motivated by several metric learning methods [44, 53], we introduce a parameter-free classification method to solve the above issue.
@@
where $\mathbf{u}_{k}$ is the centroid of class $c_{k}$, which is the average output of the predicted network $\Omega$ over samples belonging to class $c_{k}$, as in [40]. <,>$\cdot$,$\cdot$>< ⋅ , ⋅ > denotes the cosine similarity of two vectors, and $\tau$ is a learnable temperature parameter. $f(x_{j};\mathcal{F}(D_{i},\Omega;H_{\theta}))$ is the output of the target network $\Omega$ based on the input $x_{j}$. Then, the parameter-free classification loss can be defined as:
@@
To further improve the performance of the model, we add a full classification head $\mathcal{Q}_{\varphi}$ parameterized by $\varphi$ as an auxiliary task for training our hypernetwork, motivated by [53]. The classification head aims to map the output of the target network $\Omega$ to probabilities of the whole classes $\mathcal{C}^{tr}$ from $\mathcal{D}^{train}$. The full classification loss is defined as:
@@
Moreover, to make parameter-free based prediction and full classification based prediction consistent, which is motivated by [54, 55], we introduce a Kullback-Leibler Divergence loss to encourage their predicted probabilities to be matched:
@@
where $KL$ is the Kullback-Leibler Divergence. $p(y|x_{j})$ and $q(y|x_{j})$ are the predicted probabilities of $x_{j}$ of parameter-free based and full classification based methods, respectively. The probabilities of corresponding classes in $p(y|x_{j})$ are padding with zero to match the dimension of $q(y|x_{j})$.
@@
where $D_{i}$ can be regarded as a task similar to that in meta-learning. By minimizing (11), our hypernetwork can be well trained. For an unseen data in testing, we utilize our hypernetwork to directly predict its network parameters, and use the parameter-free based method for classification.
We provide the training procedure of our PudNet as listed in Algorithm 1. For each training dataset $D_{i}\in\mathcal{D}^{train}$, we first derive the sketch $\mathbf{s}_{i}$ of dataset $D_{i}$ and set the initial hidden state $\mathbf{d}_{0}^{(i)}=\mathbf{s}_{i}$ in AHRU. Then, we predict the parameters of each layer in the target network $\Omega$. Finally, we optimize the the learnable parameters $\theta$, $\varphi$ by the the overall multi-task loss $\mathcal{L}_{train}$. We also present the inference process of our PudNet for an unseen dataset, as listed in Algorithm 2.
@@ 6               Initialize the hidden state $\mathbf{d}_{0}^{(i)}=\mathbf{s}_{i}$;
8                   Predict the parameters $w_{t}$ of the $t$-th layer by weight generator;
9                   Fix the predicted parameters $w_{t}$ to $\Omega$;
@@ 12                   Compute the parameter-free loss $\mathcal{L}_{1}$ with a batch size of $b$;
Input: An unseen dataset $D_{test}\in\mathcal{D}^{test}$, the learned PudNet $H_{\theta}$, target network architecture $\Omega$.
@@
2         Predict the parameters $w_{t}$ of the $t$-th layer by the PudNet $H_{\theta}$;
3         Fix the predicted parameters $w_{t}$ to $\Omega$;
@@
5  Conduct inference for dataset $D_{test}$ with parameterized network $\Omega$ ;
Output: The prediction results for dataset $D_{test}$.
@@ III-F Theoretical Analysis of the Hyper-Mapping
We provide a theoretical analysis of how the model learns hyper-mapping between datasets and their corresponding network parameters, which offers interpretability.
Recall that for each image dataset $D_{i}\in\mathcal{D}^{train}$, the learning of the hyper-mapping between dataset $D_{i}$ and their corresponding network parameters $\mathcal{M}_{i}$ can be formulated as: leveraging a hyper-network $H_{\theta}$ to produce an optimal model $\Omega_{\mathcal{M}_{i}^{*}}$ with parameters $\mathcal{M}_{i}^{*}=H_{\theta}(D_{i})$ via optimizing $\mathop{\arg\min}\limits_{\theta}\sum_{D_{i}\in\mathcal{D}^{train}}\mathcal{L}(\Omega_{\mathcal{M}_{i}^{*}},D_{i})$. Here we denote $D_{i}$ as $d$, and we can learn the hypernetwork $H_{\theta}$ by solving an optimization problem:
@@
It is generally feasible to solve $H_{\theta}$. We use $h$ to represent $H_{\theta}$ and instead optimize the problem with the form $h^{*}:=\mathop{\arg\min}_{h}\mathcal{L}(h(d),d)$. For $h$, conditioning $\mathcal{L}(h^{*}(d),d)$ is an extremum for any infinitesimal $\delta d$, we can deduce $h^{*}$ using the ordinary differential equation:
@@
We step further by utilizing the product rule and chain rule to compute the derivative $\frac{\partial h^{*}}{\partial d}$ and substitute $\frac{\partial\mathcal{L}}{\partial h}=\frac{\partial\mathcal{L}}{\partial y}\frac{\partial y}{\partial h}$ into the equation, where $y$ is the output of the ConvNet model. Given these, $\frac{\partial^{2}\mathcal{L}}{\partial h^{2}}$ is the Hessian matrix of $\mathcal{L}$ with respect to $h$ as:
@@
Suppose we have found a solution $h^{*}$ of the optimization problem $\mathop{\arg\min}_{h^{*}}\mathcal{L}(h^{*},d_{0})$ for the dataset $d=d_{0}$. We want to use this solution to ”track” the local minimum of the loss $\mathcal{L}$ as we move from $d_{0}$ to another dataset $d=d_{1}$ in a small vicinity of $d_{0}$, where the loss $\mathcal{L}$ remains convex and the Hessian of the loss $\mathcal{L}$ is not singular.
To do this, we can use an ordinary differential equation (ODE) approach to integrate the solution with respect to a path $\hat{d}$ that connects the two datasets $d_{0}$ and $d_{1}$. To track the local minimum of the loss function as we move from $d_{0}$ to $d_{1}$ along the path $\hat{d}$, we need to integrate the above path $\hat{d}$ with $\partial h^{*}/\partial d$ as the derivative, where $d_{0}=\hat{d}(0)$ and $d_{1}=\hat{d}(1)$. By solving this ODE numerically, we obtain the solution that correspond to the local minimum of the loss $\mathcal{L}$:
@@
So far, the above equation gives us the way to get the empirical solution of $H_{\theta}$ directly with respect to the proposed PudNet. These aforementioned concepts provide a theoretical basis and the interpretability of hyper-mapping learning between datasets and their corresponding network parameters.

### v3

@@ III-D Context-Aware Parameter Prediction
After obtaining the sketches for all training datasets, we will feed them into the context-aware parameter prediction module, as shown in Fig. 2. In the following, we will introduce this module in detail.
III-D1 Capturing Contextual Parameter Relations via Adaptive Hyper-Recurrent Units
First, since we aim to establish a hypernetwork that can capture the hyper-mapping between datasets and their corresponding network parameters, it is essential for the hypernetwork to be adaptable to different datasets. This adaptability allows it to predict dataset-specific parameters effectively. Second, as the input of a neural network would sequentially pass forward its layers, the parameters of different layers are not independent. Neglecting the contextual relations among parameters from different layers may result in sub-optimal solutions.
@@ III-D1 Capturing Contextual Parameter Relations via Adaptive Hyper-Recurrent Units
In light of this, we construct Adaptive Hyper-Recurrent Units (AHRU), specifically designed for parameter prediction of ConvNets. AHRU enhances the capabilities of the gated recurrent unit by adapting to different datasets, while also serving as the hyper-recurrent unit to capture context-aware parameter relations during the parameter generation process. As shown in Fig. 3, we utilize the dataset sketch embeddings $\mathbf{s}_{i}$ as the input to the hyper-recurrent unit:
$\mathbf{d}_{0}^{(i)}=\mathbf{s}_{i},\quad\mathbf{s}_{i}\in\{\mathbf{s}_{i}\}_{i=1}^{\mathcal{N}},$ (4)
where $\mathbf{d}_{0}^{(i)}$ represents the initial hidden state of AHRU. In this way, we can integrate dataset-specific information into parameter prediction.
Then, we predict parameter representation $\mathbf{a}_{t}^{(i)}$ of $t$-th layer in $\Omega$ for the $i$-th dataset as follows:
@@
where $\mathbf{d}_{t-1}^{(i)}$ serves as the hidden state in AHRU for conveying both dataset-related context information specific to the $i$-th dataset and parameter information from previous layers. $\mathbf{W}_{r},\mathbf{W}_{z},\mathbf{W}_{h},\mathbf{W}_{o}$ are learnable parameters shared across different datasets. In AHRU, the reset gate $\mathbf{r}_{t}^{(i)}$ decides how much data information in hidden state $\mathbf{d}_{t-1}^{(i)}$ needs to be reset. The new memory $\tilde{\mathbf{d}}_{t}^{(i)}$ absorbs information from both $\mathbf{d}_{t-1}^{(i)}$ and $\mathbf{a}_{t-1}^{(i)}$. The update gate $\mathbf{z}_{t}$ regulates how much information in $\tilde{\mathbf{d}}_{t}^{(i)}$ to update and how much information in $\mathbf{d}_{t-1}^{(i)}$ to forget. During this process, the hidden state $\mathbf{d}_{t-1}^{(i)}$, along with the parameter representation $\mathbf{a}_{t-1}^{(i)}$ in the previous layer, are utilized to produce $\mathbf{a}_{t}^{(i)}$. By doing so, $\mathbf{a}_{t}^{(i)}$ can effectively encompass the dependency information between different layers while also adapting to the $i$-th dataset. The derived parameter representation $\mathbf{a}_{t}^{(i)}$ is then taken as input to the weight generator for predicting the parameters of the $t$-th layer of the target network $\Omega$.
It is worth noting that AHRU is designed based on the GRU-based architecture [52] but differs from the traditional GRU. Our AHRU is specifically designed to adapt to various datasets based on dataset sketches, while also capturing dependencies among parameters across different layers. Specifically, conventional GRU utilizes the randomly initialized hidden state to carry the long-term memory. In each recurrent step, it needs to receive external input (e.g., word) to update the hidden state and capture the contextual information. Instead, we employ each dataset sketch embedding as input and follow a self-loop structure to incorporate the predicted structure parameter representation from the previous layer. This approach enables us to leverage dataset-specific information and meanwhile allows the information of shallower layer parameters to assist in predicting parameters in deeper layer. By utilizing AHRU, our PudNet possesses the ability to adapt to various datasets while effectively capturing dependencies among parameters across different layers.
@@ III-D2 Initial Residual Connection
To ensure that the final context-aware output retains a portion of the initial dataset information, we present an initial residual connection between the dataset sketch embedding $\mathbf{s}_{i}$ and $\mathbf{a}_{t}$ as:
$\hat{\mathbf{a}}_{t}^{(i)}=\mathbf{a}_{t}^{(i)}\times(1-\eta)+\mathbf{s}_{i}\times\eta,\quad\hat{\mathbf{a}}_{t}^{(i)}\in\mathbb{R}^{m},$ (5)
where $\eta$ is the hyperparameter. After obtaining $\hat{\mathbf{a}}_{t}^{(i)}$, we input $\hat{\mathbf{a}}_{t}^{(i)}$ into the weight generator to generate parameters of the $t$-th layer of the target network $\Omega$.
@@ III-D3 Weight Generator
Since the target network $\Omega$ usually has varying sizes and dimensions across different layers, we construct a weight generator $g_{\psi_{t}}$ for each layer $t$ to transform the fixed-dimensional $\hat{\mathbf{a}}_{t}^{(i)}$ into a network parameter tensor $\mathbf{w}_{t}^{(i)}$ with variable dimensions. Here $g_{\psi_{t}}$ denotes the weight generator for the $t$-th layer, $\psi_{t}$ is the learnable parameters of $g_{\psi_{t}}$, and $\mathbf{w}_{t}^{(i)}$ is the predicted parameter of the $t$-th layer in $\Omega$ for the $i$-th dataset. We can derive the predicted parameter of the $t$-th layer as:
@@
where $g_{\psi_{t}}$ consists of one linear layer and two $1\times 1$ convolutional layers. Fig. 4 shows the architecture of the weight generator. The weight generator takes the $m$-dimensional vector $\hat{\mathbf{a}}_{t}^{(i)}$ as input and outputs a tensor with dimensions $C_{t}^{out}\times C_{t}^{in}\times k_{t}^{2}$. This tensor is the parameters generated for the $t$-th convolutional layer in $\Omega$. In a typical deep convolutional network, the model parameters are in the kernels of convolutional layers. We suppose that the $t$-th convolutional layer contains $C_{t}^{out}\times C_{t}^{in}$ kernels and each kernel has dimensions $k_{t}\times k_{t}$. To generate weights for the convolutional layers, we introduce a weight generator which consists of two parts. Firstly, we take the context-aware dataset embedding as input to generate a weight embedding of size $d=pk_{t}k_{t}$. Secondly, the weight embedding is reshaped to $(p,k_{t}k_{t})$ to get $p$ flattened kernels, where $p$ is a hyper-parameter. These $p$ flattened kernels are passed through $1\times 1$ convolutional layers and an activation layer to obtain the desired $C_{t}^{out}C_{t}^{in}$ flattened kernels. Subsequently, the matrix with dimensions $C_{t}^{out}C_{t}^{in}\times k_{t}k_{t}$ is reshaped into a tensor with dimensions $C_{t}^{out}\times C_{t}^{in}\times k_{t}\times k_{t}$ as the predicted parameter $\mathbf{w}_{t}^{(i)}$.
@@ III-E Optimization of Our Framework
In this section, we present the optimization process for our PudNet. Unlike traditional classification tasks where the training and testing data share identical label spaces, our task involves label spaces that can differ and may not overlap between training and testing datasets. Thus, training a classification head on the training data cannot be directly used to predict labels of the testing data. Motivated by several metric learning methods [44, 53], we introduce a parameter-free classification method to address the above challenge.
@@
where the centroid of class $c_{k}$, denoted by $\mathbf{u}_{k}$, represents the average output of the predicted network $\Omega$ over samples belonging to class $c_{k}$, as described in [40]. <,>$\cdot$,$\cdot$>< ⋅ , ⋅ > denotes the cosine similarity between two vectors, and $\tau$ is a learnable temperature parameter. $f(x_{j};\mathcal{F}(D_{i},\Omega;H_{\theta}))$ is the output of the target network $\Omega$ based on the input $x_{j}$. Then, the parameter-free classification loss can be defined as follows:
@@
To further improve the performance of the model, we incorporate a full classification head $\mathcal{Q}_{\varphi}$, parameterized by $\varphi$, as an auxiliary task during the training of our hypernetwork, motivated by [53]. The full classification head aims to map the output of the target network $\Omega$ into probabilities of whole classes $\mathcal{C}^{tr}$ in the dataset $\mathcal{D}^{train}$. The full classification loss is defined as:
@@
Furthermore, to ensure the consistency between parameter-free classification prediction and full classification prediction, as motivated by [54, 55], we introduce a Kullback-Leibler Divergence loss to encourage the alignment of their predicted probabilities:
@@
where $KL$ is the Kullback-Leibler Divergence. $p(y|x_{j})$ and $q(y|x_{j})$ denote the predicted probabilities of $x_{j}$ obtained from the parameter-free based and full classification based methods, respectively. To ensure dimensional compatibility, the probabilities of the corresponding classes in $p(y|x_{j})$ are padded with zeros to match the dimension of $q(y|x_{j})$.
@@
where $D_{i}$ can be regarded as a task analogous to it in meta-learning. By minimizing (11), our hypernetwork can be effectively trained. During testing, for unseen data, we employ our hypernetwork to directly predict its network parameters and utilize the parameter-free classification method for classification.
We provide the training procedure of our PudNet in Algorithm 1. For each training dataset $D_{i}\in\mathcal{D}^{train}$, we first derive the sketch $\mathbf{s}_{i}$ of dataset $D_{i}$ and initialize the hidden state as $\mathbf{d}_{0}^{(i)}=\mathbf{s}_{i}$ in AHRU. Then, we predict the parameters of each layer in the target network $\Omega$. Finally, we optimize the learnable parameters $\theta$ and $\varphi$ using the overall multi-task loss $\mathcal{L}_{train}$. We also present the inference process of our PudNet for an unseen dataset in Algorithm 2.
@@ 6               Initialize the hidden state $\mathbf{d}_{0}^{(i)}=\mathbf{s}_{i}$;
8                   Predict the parameters $w_{t}^{(i)}$ of the $t$-th layer by weight generator;
9                   Fix the predicted parameters $w_{t}^{(i)}$ to $\Omega$;
@@ 12                   Compute the parameter-free loss $\mathcal{L}_{1}$ with a batch size of $b$;
Input: An unseen dataset $D_{i}\in\mathcal{D}^{test}$, the learned PudNet $H_{\theta}$, target network architecture $\Omega$.
@@
2         Predict the parameters $w_{t}^{(i)}$ of the $t$-th layer by the PudNet $H_{\theta}$;
3         Fix the predicted parameters $w_{t}^{(i)}$ to $\Omega$;
@@
5  Conduct inference for dataset $D_{i}$ with parameterized network $\Omega$ ;
Output: The prediction results for dataset $D_{i}$.
@@ III-F Theoretical Analysis of the Hyper-Mapping
We provide a theoretical analysis elucidating how the model learns the hyper-mapping between datasets and their corresponding network parameters, which offers interpretability.
Recall that for each image dataset $D_{i}\in\mathcal{D}^{train}$, the learning of the hyper-mapping between dataset $D_{i}$ and their corresponding network parameters $\mathcal{M}_{i}$ can be formulated as: leveraging a hyper-network $H_{\theta}$ to produce an optimal model $\Omega_{\mathcal{M}_{i}^{*}}$ with parameters $\mathcal{M}_{i}^{*}=H_{\theta}(D_{i})$ via optimizing $\mathop{\arg\min}\limits_{\theta}\sum_{D_{i}\in\mathcal{D}^{train}}\mathcal{L}(\Omega_{\mathcal{M}_{i}^{*}},D_{i})$. To write conveniently, we denote $D_{i}$ as $d$. We can learn the hypernetwork $H_{\theta}$ by solving an optimization problem:
@@
Solving $H_{\theta}$ is generally feasible. We denote $H_{\theta}$ as $h$ and optimize the problem in the form $h^{*}:=\mathop{\arg\min}_{h}\mathcal{L}(h(d),d)$. For $h$, conditioning $\mathcal{L}(h^{*}(d),d)$ is an extremum for any infinitesimal $\delta d$, we can deduce $h^{*}$ using the ordinary differential equation:
@@
We proceed by employing the product rule and chain rule to calculate the derivative $\frac{\partial h^{*}}{\partial d}$. We substitute $\frac{\partial\mathcal{L}}{\partial h}=\frac{\partial\mathcal{L}}{\partial y}\frac{\partial y}{\partial h}$ into the equation, where $y$ represents the output of the ConvNet model. Given these, $\frac{\partial^{2}\mathcal{L}}{\partial h^{2}}$ is the Hessian matrix of $\mathcal{L}$ with respect to $h$ as:
@@
Suppose we have found a solution $h^{*}$ for the optimization problem $\mathop{\arg\min}_{h^{*}}\mathcal{L}(h^{*},d_{0})$ concerning the dataset $d=d_{0}$. We intend to utilize this solution to ”track” the local minimum of the loss $\mathcal{L}$ as we move from $d_{0}$ to another dataset $d=d_{1}$ within a small vicinity of $d_{0}$, where the loss $\mathcal{L}$ remains convex and the Hessian of the loss $\mathcal{L}$ is non-singular.
To accomplish this, we can employ an ordinary differential equation (ODE) approach to integrate the solution along a path $\hat{d}$ that connects the two datasets $d_{0}$ and $d_{1}$. To track the local minimum of the loss function when transitioning from $d_{0}$ to $d_{1}$ along the path $\hat{d}$, we need to numerically integrate the above path $\hat{d}$ with $\partial h^{*}/\partial d$ as the derivative, where $d_{0}=\hat{d}(0)$ and $d_{1}=\hat{d}(1)$. By solving this ODE numerically, we obtain the solution corresponding to the local minimum of the loss $\mathcal{L}$:
@@
The aforementioned equation provides us the way to obtain the empirical solution of $H_{\theta}$ in relation to the proposed PudNet. These concepts establish a theoretical basis and enhance the interpretability of hyper-mapping learning between datasets and their corresponding network parameters.